# Predicting from Strings: Language Model Embeddings for Bayesian Optimization

### v2

@@ 4.2 Ablations
Figure 7: (,,,)($\downarrow$,$\downarrow$,$\uparrow$,$\downarrow$)( ↓ , ↓ , ↑ , ↓ ) are better, respectively. Number of historical context points vs predictive metrics on unseen points over unseen BBOB function trajectories, while varying string embedder sizes. Solid line denotes mean over 10 test functions and error bars denote standard deviation.
@@ 5 Conclusion and Future Work
Our method, Embed-then-Regress, demonstrates the versatility of using string-based in-context regression for Bayesian Optimization over a variety of problems. We have shown it to obtain comparable results against industry-standard GP baselines and allow flexibility in more esoteric spaces such as permutations and combinations.
As strings are significantly more flexible representation formats of different data types, an ambitious and exciting direction is to pretrain a unified in-context regression model over multiple different domains, in order to obtain a “universal” in-context regressor. Numerous architectural improvements could also be made. For example, the aggregation method over Transformer outputs may be learned rather than predefined using average pooling as in this paper, and more broadly there may be better methods for computing fixed-length embeddings from a Transformer. It may also be possible to additionally create a string-based GP by sending string embeddings as inputs to a kernel, for better guarantees of uncertainty estimates.
Further possible applications include prompt optimization (Fernando et al., 2024) and code search (Romera-Paredes et al., 2023), areas which still predominantly use zeroth-order evolutionary algorithms or even random search, which can be very sample inefficient compared to Bayesian Optimization. Additionally, outside of blackbox optimzation problems which are stateless with respect to inputs, it is worth investigating whether such methods are applicable for process-based reward modelling (Lightman et al., 2024) and tree search-based approaches (Yao et al., 2023) for stateful environments in language modelling.
@@ Appendix A Model Details
  • Training: Effective batch size of 16, learning rate of $5\times 10^{-4}$, weight decay of $10^{-5}$, gradient clipping of 0.5, using the AdamW optimizer. A fixed number $T\geq 100$ total trials were always placed in the context window, with the number of history trials $t^{\prime}$ sampled between $[10,T-10]$ and the rest were target points for loss computations.
@@
For combinatorial tasks, Regularized Evolution used a population size of 50, and a tournament size of 7population $7\approx\sqrt{\text{population size}}$, as prescribed in (Real et al., 2019). When it was augmented by the acquisition, we chose the highest scoring proposal out of 5 samples as the final candidate for evaluation. The acquisition-based ranking began at trial 20.
@@ Appendix B Benchmarking
[…]
@@ B.1 BBOB
  • Shifting: Uniformly samples a shift $c\in\mathbb{R}^{\text{dim}}$, and transform $f(x)$ into $f(x-c)$.
@@
  • Discretization: Each parameter is randomly chosen to remain continuous, or either a DISCRETE or CATEGORICAL parameter, whose feasible points are selected over a uniform grid between the original bounds $[-5,5]$, with the number of feasible points uniformly selected from 2 to 16.
@@ B.2 Combinatorial
Permutation: Let $x$ be a permutation of $[n]=\{1,2,\ldots,n\}$ where $x^{(i)}$ denotes the permutation index at position $i$.

### v3

@@ 4.2 Ablations
Figure 7: $(\downarrow,\downarrow,\uparrow,\downarrow)$ are better, respectively. Number of historical context points vs predictive metrics on unseen points over unseen BBOB function trajectories, while varying string embedder sizes. Solid line denotes mean over 10 test functions and error bars denote standard deviation.
@@ 5 Conclusion and Future Work
We investigated the use of string-based ICL regression for Bayesian Optimization over a variety of synthetic, combinatorial, and real-world optimization problems, and shown it to obtain comparable results against industry-standard GP baselines, yet allow flexibility in esoteric spaces such as permutations and combinations. We leave architectural improvements for future work - for example, the aggregation method over Transformer outputs may be learned rather than predefined using average pooling as in this paper, and more broadly there may be better methods for computing fixed-length embeddings from a Transformer. Further investigation into different regression bodies could lead to e.g. string-based GPs using kernels over string embeddings for better guarantees of uncertainty estimates.
An ambitious and exciting direction is to pretrain a unified in-context regression model broadly over multiple different domains including prompt optimization (Fernando et al., 2024) and code search (Romera-Paredes et al., 2023), in order to obtain a “universal” in-context regressor which can speed up search over evolutionary algorithms, as strings are significantly more flexible representation formats of different data types. Additionally, outside of optimization problems which are stateless with respect to inputs, it is worth investigating whether such methods are applicable for LLM reasoning as reward models to assist tree search-based approaches (Yao et al., 2023) for stateful environments in language modeling.
@@ Appendix A Model Details
  • Training: Effective batch size of 16, learning rate of $5\times 10^{-4}$, weight decay of $10^{-5}$, gradient clipping of 0.5, using the AdamW optimizer. A fixed number $T\geq 100$ total trials were always placed in the context window, with the number of history trials $t^{\prime}$ sampled between [10,T10][10,T-10] and the rest were target points for loss computations.
@@
For combinatorial tasks, Regularized Evolution used a population size of 50, and a tournament size of $7\approx\sqrt{\text{population size}}$, as prescribed in (Real et al., 2019). When it was augmented by the acquisition, we chose the highest scoring proposal out of 5 samples as the final candidate for evaluation. The acquisition-based ranking began at trial 20.
@@ Appendix B Benchmarking
For every algorithm and objective pair, we run 20 seeds and plot the best-so-far mean with $0.5*\text{(standard deviation)}$ as error bars.
@@ B.1 BBOB
  • Shifting: Uniformly samples a shift $c\in\mathbb{R}^{\text{dim}}$, and transform $f(x)$ into f(xc)f(x-c).
@@
  • Discretization: Each parameter is randomly chosen to remain continuous, or either a DISCRETE or CATEGORICAL parameter, whose feasible points are selected over a uniform grid between the original bounds [5,5][-5,5], with the number of feasible points uniformly selected from 2 to 16.
@@ B.2 Combinatorial
Permutation: Let $x$ be a permutation of $n]=\{1,2,\ldots,n\}$ where $x^{(i)}$ denotes the permutation index at position $i$.